The distance between the objective and eyepeices lenses in a microscope is 20 cm. The objective lens has a focal length of 5.0mm. What eye piece focal lengthe will give the microscope an overall angular magnification of 250?
Answer #1

Knowing that a standard microscope has a tube length and hence objective image distance of 160 mm and that the objective object distance is essentially the objective focal length, we can determine the magnification of the objective lens. Knowing the magnification of the objective and the overall magnification we can determine the magnification of the eyepiece.

The magnification of the objective is obtained by

m0=-(S0')/(S0)=-L/f0

m0=-160mm/5.0mm

m0=-32

Knowing the overall magnification we can determine the magnification of the eyepiece by

M = m0me

me=M/m0

me=-250/-32=7.81

me=7.8